# Large Deviations Principle for Isoperimetry and Its Equivalence to Nonlinear Log-Sobolev Inequalities

### v3

@@ 1.3. Related Works
In the one-dimensional setting, Bobkov [11] established that half-lines are extremal sets in the isoperimetric inequality for log-concave probability measures. Extending to higher-dimensional structures, Bobkov and Houdré [15] further explored the isoperimetric constant and log-Sobolev inequalities in product spaces, laying a foundation for multi-dimensional isoperimetric analysis. The optimality of standard half-spaces— a core topic in isoperimetry— has been systematically studied by multiple scholars: Sudakov and Tsirel’son [61] initiated early investigations, Bobkov [11] deepened this topic by addressing the isoperimetric problem for log-concave product measures on $\mathbb{R}^{n}$ equipped with the uniform distance, and explicitly provided necessary and sufficient conditions for standard half-spaces to be extremal, and Barthe [9] investigated the extremal properties of central half-spaces (with measure $1/2$) for product probability measures on Riemannian manifolds equipped with the geodesic distance. It should be noting that the isoperimetric problem under the geodesic distance and the one under the uniform distance are significantly different. When equipped with the uniform distance, the asymptotics of the isoperimetric profile has been characterized by Bobkov in [12]. However, for the geodesic distance as in our setup, the asymptotics is still widely unknown except for special cases, e.g., the Gaussian setting. Roberto [59] offers a critical survey for a comprehensive overview of isoperimetry in product spaces. Furthermore, the isoperimery for small sets was already investigated in the literature; see, e.g., [41] for the continuous setting and [56, 53] for the discrete setting. In the discrete setting, the isoperimery for small sets is known as the small set expansion and has important applications in complexity theory [56]. It should be noted that characterizing the isoperimetric profile remains widely open, and previously so does its asymptotics.
@@
Ahlswede and Zhang [3] and Yu [68] contributed to the study of isoperimetry with thick boundaries for exponentially small sets. This isoperimetric variant refines the concentration of measure phenomenon by relaxing the set size requirement to allow arbitrarily exponentially small measures, yet it behaves fundamentally differently from—and is far more intricate than—concentration of measure. This is because, in contrast to the latter, the isoperimetric variant admits no dimension-independent bounds. Ahlswede and Zhang [3] focused on finite spaces for which a key tool called the inherently typical subset lemma was developed by them. Yu [68] generalized their result to Polish spaces (complete, separable metric spaces), expanding the problem’s applicability, especially to the setting considered in this paper (i.e., the isoperimetric problem on a manifold with density), which plays a core role in our proof in the present paper.
@@
Ledoux [37, 39] proposed a semigroup approach to derive isoperimetric inequalities from hypercontractivity (which is equivalent to log-Sobolev inequalities). The semigroup approach can be used to establish isoperimetric inequalities for the whole range of volumes, but the constant obtained via this method is in general suboptimal. In contrast, the constant in our work is asymptotically optimal. Otto and Villani [54] established that log-Sobolev inequalities imply transport-entropy inequalities, bridging log-Sobolev theory with concentration of measure. Building on this, Ledoux [40] (alongside Milman [45]) used a semigroup approach to derive isoperimetric inequalities from concentration inequalities (equivalently, transport-entropy inequalities). As a nonlinear log-Sobolev inequality for the standard heat semigroups (with Lebesgue measures as stationary measures), the entropic isoperimetric inequality was investigated in [20, 24] which is in fact equivalent to the standard log-Sobolev inequality for Ornstein–Uhlenbeck semigroups (with Gaussian measures as stationary measures). Nonlinear log-Sobolev inequalities for general heat (diffusion) semigroups were investigated by F. Y. Wang [66], Bakry, Gentil, and Ledoux [7] (called the entropy-energy inequality therein), as well as Polyanskiy and Samorodnitsky [55]. Specifically, F. Y. Wang [66] showed that nonlinear log-Sobolev inequalities imply nonlinear transport-entropy inequalities, extending Otto and Villani’s linear version to the nonlinear setting. Bakry, Gentil, and Ledoux [7] and Polyanskiy and Samorodnitsky [55] applied nonlinear log-Sobolev inequalities to strengthen hypercontractive inequalities. Though nonlinear Sobolev inequalities have been studied in the literature not directly to address asymptotic isoperimetry, an exact equivalence between these two concepts does exist—one we establish in the present work. Furthermore, the equivalence between isoperimetry and Sobolev-type inequalities was investigated in [16].
@@ 1.4. Proof Idea and Paper Organization
Ledoux’s semigroup method, as developed in [37, 39], seems insufficient for establishing Theorem 1. This limitation stems from the method’s core strategy: it reduces the isoperimetric problem to the more general noise stability problem. However, obtaining a sufficiently precise solution to the former demands an exact (or sufficiently precise, not merely asymptotic) solution to the latter—an outcome that remains elusive in general. While asymptotically optimal solutions for the noise stability problem are well-established [53, 34, 69, 67], a complete exact resolution has yet to be achieved. Accordingly, a novel approach is required to advance the proof of Theorem 1.
@@
The heuristic reasoning behind our idea is as follows. Given an arbitrary set in a high-dimensional space, its (weighted) surface area does not admit any explicit formula and thus, is difficult to analyze. How can we convert the surface area to a simple formula? A possible way is to tailor the set to a new one that is “regular” in a certain sense, so that the surface area of the resulting set admits a simple formula. However, this is difficult to realize in a direct way, since the surface area is very sensitive to the shape of the set, and thus any subtle change of the set could lead to drastic change in surface area.
@@
Our proof idea of Theorem 1 is as follows, which is new and totally different from the semigroup approach. Firstly, using a generalized Heintze-Karcher theorem [49] from geometric measure theory, we convert the standard isoperimetric problem into a variant version in which the “perimeter” is measured by the size of a thick boundary of the set. Secondly, using the inherently typical subset lemma [2] and its consequence [68] from information theory, we tailor the set into a “regular” one whose thick boundary admits a simple formula. The formula we obtain is expressed in terms of nonlinear transport-entropy inequality. The feasibility of this step is due to the fact that the size of thick boundary is robust to any slight change of the set. Lastly, using nonlinear transport inequality [66, 65] from optimal transport, we simplify the resulting formula and express it in terms of nonlinear log-Sobolev inequality, leading to the equivalence between the standard isoperimetric problem and the nonlinear log-Sobolev inequality.
@@
Our specific strategy is illustrated in Fig. 1.1, which is described in the order opposite to the one given above. Specifically, using tools from optimal transport, in Section 2 we first deduce a specific form of nonlinear transport inequality from the nonlinear log-Sobolev inequality. By tools from information theory, we show in Section 3 that the nonlinear transport inequality in turn implies a variant isoperimetric inequality—the isoperimetric inequality with thick boundaries. Lastly, by tools from geometric measure theory, we show in Section 4 that the variant isoperimetric inequality implies the original isoperimetric inequality. Putting everything together, in Section 5 we deduce that the nonlinear log-Sobolev inequality implies the isoperimetric inequality. As for the other direction, we provide two proofs in Section 6: one is based on information theory and optimal transport, and the other is based on geometric measure theory.
We prove Theorem 2, Theorem 4, Theorem 5, Proposition 2, and Proposition 3 respectively in Appendices A–E.
@@ 2.1. From Log-Sobolev to Transport
In this section, we connect log-Sobolev inequality and transport-entropy inequality. Recall that $\mathcal{M}$ is a Riemannian manifold, equipped with a Borel reference measure $\nu=e^{-V}\mathrm{vol}\in P_{2}^{\mathrm{ac}}(\mathcal{M})$ with $V\in C^{2}(\mathcal{M})$. Let $\mu$ be a probability measure on $\mathcal{M}$, whose density w.r.t. $\nu$ is $\rho$. Recall that the entropy and the Fisher information are respectively
$D(\mu\|\nu)$ $:=\int\rho\ln\rho{\text{\rm d}}\nu,$
$I(\mu\|\nu)$ $:=\int\frac{|\nabla\rho|^{2}}{\rho}{\text{\rm d}}\nu.$
@@
It is well known that Sobolev-type inequalities (including log-Sobolev inequalities) imply Talagrand’s transport inequalities; this point was shown by Otto and Villani [54]. A nonlinear analogue of this result was given by F. Y. Wang [66, Theorem 2.1] which is proven by extending Otto and Villani’s idea. Specifically, F. Y. Wang showed that if
(2.1) $\theta(D(\lambda\|\nu))\leq I(\lambda\|\nu),\;\forall\lambda\in P_{2}^{\mathrm{ac}}(\mathcal{M})$
for some function $\theta$, then
@@
(2.2) $\mathsf{W}_{2}\left(\mu,\nu\right)\leq\int_{0}^{D(\mu\|\nu)}\frac{1}{\sqrt{\theta(r)}}{\text{\rm d}}r,\;\forall\mu\in P_{2}^{\mathrm{ac}}(\mathcal{M}).$
In the present paper, we require a variant of F. Y. Wang’s result.
@@
Assume that (2.1) holds for some function $\theta$. Then, for any $\mu\in P_{2}^{\mathrm{ac}}(\mathcal{M})$ and any $0\leq s\leq D(\mu\|\nu)$, there is a probability measure $\pi\in P_{2}^{\mathrm{ac}}(\mathcal{M})$ such that $D(\pi\|\nu)=s$ and
(2.3) $\mathsf{W}_{2}\left(\mu,\pi\right)\leq\int_{s}^{D(\mu\|\nu)}\frac{1}{\sqrt{\theta(r)}}{\text{\rm d}}r.$
@@
Define a Laplace-type operator $L=\Delta-\nabla V\cdot\nabla$. Let $\mu\in P_{2}^{\mathrm{ac}}(\mathcal{M})$ be a probability measure with density $h=\frac{{\text{\rm d}}\mu}{{\text{\rm d}}\nu}$ such that
[…]
The heat equation $\partial_{t}\rho=L\rho$ with initial data $\rho_{0}=h$ admits a solution $\rho_{t}$; see [54]. Let $\mu_{t}=\rho_{t}\nu$. The heat equation has intimate relations with the entropy and the Wasserstein metric in the sense that the heat equation is the gradient flow of the entropy functional $D$ with respect to the Wasserstein metric $\mathsf{W}_{2}$.
By an approximation argument in [54], for every $\mu\in P_{2}^{\mathrm{ac}}(M)$, there is a sequence of probability measures $\mu_{k}$ such that the density $h_{k}:={\text{\rm d}}\mu_{k}/{\text{\rm d}}\nu$ for each $k$ satisfies (2.4) and moreover, as $k\to\infty$,
(2.5) $\mathsf{W}_{2}\left(\mu_{k},\mu\right)$ $\to 0,$
(2.6) $D(\mu_{k}\|\nu)$ $\to D(\mu\|\nu).$
Given any $0<s<D(\mu\|\nu)$, it holds that $0<s<D(\mu_{k}\|\nu)$ for all sufficiently large $k$. Following the proof of F. Y. Wang [66, Theorem 2.1] (taking the integration $\int_{t}^{\infty}$ for a proper $t$, instead of $\int_{0}^{\infty}$ [66, (2.5)]), it is obtained that for each such $k$, there is a probability measure $\pi_{k}$ (the heat flow at $t$ with initial data $\mu_{k}$) such that $D(\pi_{k}\|\nu)=s$ and
(2.7) $\mathsf{W}_{2}\left(\mu_{k},\pi_{k}\right)\leq\int_{s}^{D(\mu_{k}\|\nu)}\frac{1}{\sqrt{\theta(r)}}{\text{\rm d}}r.$
Note that (2.6) implies that $\mu_{k}$ belongs to a sublevel set of the relative entropy $\pi\mapsto D(\pi\|\nu)$, and so does $\pi_{k}$. Since the sublevel set of the relative entropy is weakly compact, $\pi_{k}$ converges to some $\pi^{*}$ weakly. Since $\mu$ has finite second moment, (2.5) implies that the sequence $\mu_{k}$ has bounded second moment. In this case, $\pi_{k}\to\pi^{*}$ weakly is equivalent to $\mathsf{W}_{2}\left(\pi_{k},\pi^{*}\right)\to 0$. Taking limits for (2.7) and applying the triangle inequality yield
(2.8) $\mathsf{W}_{2}\left(\mu,\pi^{*}\right)\leq\int_{s}^{D(\mu\|\nu)}\frac{1}{\sqrt{\theta(r)}}{\text{\rm d}}r.$
Moreover, $D(\pi^{*}\|\nu)=s$. Such $\pi^{*}$ is what we want. ∎

### v4

@@ 1.2. Related Works
In the one-dimensional setting, Bobkov [17] established that half-lines are extremal sets in the isoperimetric inequality for log-concave probability measures. Extending to higher-dimensional structures, Bobkov and Houdré [21] further explored the isoperimetric constant and log-Sobolev inequalities in product spaces, laying a foundation for multi-dimensional isoperimetric analysis. The optimality of standard half-spaces— a core topic in isoperimetry— has been systematically studied by multiple scholars: Sudakov and Tsirel’son [68] initiated early investigations, Bobkov [17] deepened this topic by addressing the isoperimetric problem for log-concave product measures on $\mathbb{R}^{n}$ equipped with the uniform distance, and explicitly provided necessary and sufficient conditions for standard half-spaces to be extremal, and Barthe [12] investigated the extremal properties of central half-spaces (with measure $1/2$) for product probability measures on Riemannian manifolds equipped with the geodesic distance. It should be noting that the isoperimetric problem under the geodesic distance and the one under the uniform distance are significantly different. When equipped with the uniform distance, the asymptotics of the isoperimetric profile has been characterized by Bobkov in [18]. However, for the geodesic distance as in our setup, the asymptotics is still widely unknown except for special cases, e.g., the Gaussian setting. Roberto [64] offers a critical survey for a comprehensive overview of isoperimetry in product spaces. See also other related works or survey in [14, 13, 65, 51]. Furthermore, the isoperimery for small sets was already investigated in the literature; see, e.g., [46] for the continuous setting and [61, 57] for the discrete setting. In the discrete setting, the isoperimery for small sets is known as the small set expansion and has important applications in complexity theory [61]. It should be noted that exact characterization of the isoperimetric profile remains widely open, and previously so does its asymptotics.
@@
Ahlswede and Zhang [3] and Yu [77] contributed to the study of isoperimetry with thick boundaries for exponentially small sets. This isoperimetric variant refines the concentration of measure phenomenon by relaxing the set size requirement to allow arbitrarily exponentially small measures, yet it behaves fundamentally differently from—and is far more intricate than—concentration of measure. This is because, in contrast to the latter, the isoperimetric variant admits no dimension-independent bounds. Ahlswede and Zhang [3] focused on finite spaces for which a key tool called the inherently typical subset lemma was developed by them. Yu [77] generalized their result to Polish metric spaces (complete, separable metric spaces), expanding the problem’s applicability, which plays a key role in our proof in the present paper.
@@
Ledoux [42, 44] proposed a semigroup approach to derive isoperimetric inequalities from hypercontractivity (which is equivalent to log-Sobolev inequalities). The semigroup approach can be used to establish isoperimetric inequalities for the whole range of volumes, but the constant obtained via this method is in general suboptimal. In contrast, the constant in our work is asymptotically optimal. Otto and Villani [58] established that log-Sobolev inequalities imply transport-entropy inequalities, bridging log-Sobolev theory with concentration of measure. Building on this, Ledoux [45] (alongside Milman [52]) used a semigroup approach to derive isoperimetric inequalities from concentration inequalities (equivalently, transport-entropy inequalities). As a nonlinear log-Sobolev inequality for the standard heat semigroups (with Lebesgue measures as stationary measures), the entropic isoperimetric inequality was investigated in [26, 29] which is in fact equivalent to the standard log-Sobolev inequality for Ornstein–Uhlenbeck semigroups (with Gaussian measures as stationary measures). Nonlinear log-Sobolev inequalities for general heat (diffusion) semigroups were investigated by Bakry [9] (called the entropy-energy inequality therein), F. Y. Wang [75], Bakry, Gentil, and Ledoux [10], as well as Polyanskiy and Samorodnitsky [59]. Specifically, F. Y. Wang [75] showed that nonlinear log-Sobolev inequalities imply nonlinear transport-entropy inequalities, extending Otto and Villani’s linear version to the nonlinear setting. Bakry [9], Bakry, Gentil, and Ledoux [10], and Polyanskiy and Samorodnitsky [59] applied nonlinear log-Sobolev inequalities to strengthen hypercontractive inequalities. Though nonlinear Sobolev inequalities have been studied in the literature not directly to address asymptotic isoperimetry, an exact equivalence between these two concepts does exist—one we establish in the present work. Furthermore, equivalence between isoperimetry and Sobolev-type inequalities was investigated in [22].
@@ 1.3. Proof Idea and Paper Organization
Ledoux’s semigroup method, as developed in [42, 44], seems insufficient for establishing Theorem 1. This limitation originates from the core strategy of the method: it transforms the isoperimetric problem into a “soft boundary” variant known as the noise stability problem. However, obtaining a precise solution to the former demands an exact (or sufficiently precise, not merely asymptotic) solution to the latter—an outcome that remains elusive in general. While asymptotically optimal solutions for the noise stability problem are well-established [57, 39, 78, 76], an exact resolution has yet to be achieved. Accordingly, a novel approach is required to advance the proof of Theorem 1.
@@
The heuristic reasoning behind our proof idea is as follows. Given an arbitrary set in a high-dimensional space, due to lack of structures, its (weighted) surface area does not admit any simple formulas and thus, is difficult to analyze. To resolve this difficulty, a possible way is to tailor the set to a new one that is “regular” in a certain sense, so that the surface area of the resulting set admits a simple formula. However, this is difficult to realize in a direct way, since the surface area is very sensitive to the shape of the set, and thus any subtle change of the set could lead to drastic change in surface area.
@@
Our proof idea of Statement 1 of Theorem 1 is as follows, which is new and totally different from the semigroup approach. Firstly, using a generalized Heintze-Karcher theorem [55] from geometric measure theory, we convert the standard isoperimetric problem into a variant version in which the “perimeter” is measured by the size of a thick shell of the set. Secondly, using information-theoretic tools—the inherently typical subset lemma [2] and its consequence [77]—as well as covering arguments, we tailor the set into a “regular” one the measure of whose thick shell admits a simple formula, which is expressed in terms of nonlinear transport-entropy inequality. The feasibility of this step is due to the fact that the size of thick shell is robust to any slight change of the set. Lastly, by establishing a nonlinear transport inequality analogous to [75] using gradient flow (heat flow) from optimal transport [73], we simplify the resulting formula and express it in terms of nonlinear log-Sobolev inequality, leading to the equivalence between the standard isoperimetric problem and the nonlinear log-Sobolev inequality.
@@
Our specific strategy is illustrated in Fig. 1.1, which is described in the order opposite to the one given above. Specifically, using tools from optimal transport, in Section 2 we first deduce a specific form of nonlinear transport inequality from the nonlinear log-Sobolev inequality. By tools from information theory as well as covering and heat flow arguments, we show in Section 3 that the nonlinear transport inequality in turn implies a variant isoperimetric inequality—the isoperimetric inequality with thick boundaries. Lastly, by tools from geometric measure theory, we show in Section 4 that the variant isoperimetric inequality implies the original isoperimetric inequality. Putting everything together, in Section 5 we deduce that the nonlinear log-Sobolev inequality implies the isoperimetric inequality. As for the other direction, we provide two proofs in Section 6: one is based on information theory and optimal transport, and the other is based on geometric measure theory.
We prove Statement 2 of Theorem 1, Theorem 2, Theorem 3, Theorem 4, and Theorem 5 respectively in Appendices A–E.
@@ 2. From Log-Sobolev to Transport
In this section, we connect log-Sobolev inequality and transport-entropy inequality. Let $(\mathcal{X},d,\nu)$ be a Polish (complete and separable) metric probability measure space, equipped with the $\nu$-completion of the Borel $\sigma$-algebra.
We now need some metric notions; refer to [6, 33] for details. A curve $\gamma:[0,\infty)\to\mathcal{X}$ is said to be absolutely continuous if there exists $g\in L^{1}(0,\infty)$ such that
(2.1) $d(\gamma_{s},\gamma_{t})\leq\int_{s}^{t}g(r)\,{\text{\rm d}}r,\quad\forall 0\leq s<t<\infty.$
In this case, the metric speed of $\gamma$ is well-defined for a.e. $t$ by
$|\dot{\gamma}_{t}|:=\lim_{h\to 0}\frac{d(\gamma_{t+h},\gamma_{t})}{|h|}.$
It turns out that $|\dot{\gamma}|\in L^{1}(0,1)$ and this is the minimal $L^{1}$ function for which (2.1) holds.
Given a function $f:X\to\mathbb{R}$, the local Lipschitz constant $|\nabla f|:\mathcal{X}\to[0,\infty]$ is defined by
$|\nabla f|(x):=\limsup_{y\to x}\frac{|f(x)-f(y)|}{d(x,y)}.$
Define the Cheeger energy functional $\operatorname{Ch}:L^{2}(\mathcal{X},\nu)\to[0,\infty]$ by
(2.2) $\operatorname{Ch}(f):=\inf\liminf_{n\to\infty}\frac{1}{2}\int|\nabla f_{n}|^{2}\,{\text{\rm d}}\nu,$
where the infimum is taken among all sequences $(f_{n})$ of Lipschitz functions converging to $f$ in $L^{2}(\mathcal{X},\nu)$. Then the Sobolev space $W^{1,2}(\mathcal{X},d,\nu)$ is defined as the set of $f\in L^{2}(\mathcal{X},\nu)$ such that $\operatorname{Ch}(f)<\infty$ endowed with the norm
(2.3) $\|f\|_{W^{1,2}}:=\|f\|_{L^{2}}^{2}+2\operatorname{Ch}(f).$
A function $G\in L^{2}(\mathcal{X},\nu)$ is said to be a relaxed gradient of $f\in L^{2}(\mathcal{X},\nu)$ if there exist Lipschitz functions $f_{n}\in L^{2}(\mathcal{X},\nu)$ such that:
  • $f_{n}\to f$ in $L^{2}(\mathcal{X},\nu)$ and $|\nabla f_{n}|$ weakly converge to $\tilde{G}$ in $L^{2}(\mathcal{X},\nu)$;
  • $\tilde{G}\leq G$ $\nu$-a.e. in $\mathcal{X}$.
A function $G$ is said to be the minimal relaxed gradient of $f$ if its $L^{2}(\mathcal{X},\nu)$ norm is minimal among relaxed gradients. We shall denote by $|\nabla f|_{*}$ the minimal relaxed gradient. In terms of the minimal relaxed gradient, the Cheeger energy can be equivalently expressed as
$\operatorname{Ch}(f)=\frac{1}{2}\int|\nabla f|_{*}^{2}\,{\text{\rm d}}\nu.$
For a probability measure $\mu$ defined on $\mathcal{X}$ with density $f$ w.r.t. $\nu$, the Fisher information is defined as
$I(\mu\|\nu):=\int_{\{f>0\}}\frac{|\nabla f|_{*}^{2}}{f}{\text{\rm d}}\nu,$
which reduces to the definition in when the space is a smooth Riemannian manifold. Recall the definition of the relative entropy:
$D(\mu\|\nu)$ $:=\int f\ln f{\text{\rm d}}\nu.$
We first the following equivalent formulations of the nonlinear log-Sobolev inequality, which is a nonlinear analogue of the linear version given in [33, Proposition 1]. The proof is almost same as the linear version, and omitted here.
Proposition 2 (Equivalent formulations of the nonlinear log Sobolev inequality [33]).
Let $(\mathcal{X},d,\nu)$ be a Polish metric measure space and let $\theta$ be a continuous function. Then the following are equivalent:
  • For any Lipschitz function $f:\mathcal{X}\to[0,\infty)$ with $\int f\,{\text{\rm d}}\nu=1$, it holds
    $\theta(\int f\ln f\,{\text{\rm d}}\nu)\leq\int_{\{f>0\}}\frac{|\nabla f|^{2}}{f}\,{\text{\rm d}}\nu.$
  • For any non-negative $f\in W^{1,2}(\mathcal{X},d,\nu)$ with $\int f\,{\text{\rm d}}\nu=1$, it holds
    $\theta(\int f\ln f\,{\text{\rm d}}\nu)\leq\int_{\{f>0\}}\frac{|\nabla f|_{*}^{2}}{f}\,{\text{\rm d}}\nu.$
Given $f\in W^{1,2}(\mathcal{X},d,\nu)$, we write $\partial^{-}\mathrm{Ch}(f)\subset L^{2}(\mathcal{X},\nu)$ for the subdifferential at $f$ of the restriction to $L^{2}(\mathcal{X},\nu)$ of Cheeger’s energy, namely, $\xi\in\partial^{-}\mathrm{Ch}(f)$ if and only if
$\mathrm{Ch}(g)\geq\mathrm{Ch}(f)+\int\xi(g-f)\,\mathrm{d}\nu\quad\forall g\in L^{2}(\mathcal{X},\nu).$
Notice that the functional $\operatorname{Ch}:L^{2}(\mathcal{X},\nu)\to[0,\infty]$ is convex, lower semicontinuous and with dense domain, so that the standard theory of gradient flows on Hilbert space applies. This means that [7, Theorem 2.13] for any $f\in L^{2}(\mathcal{X},\nu)$ there exists a unique locally absolutely continuous curve $(f_{t})\subset L^{2}(\mathcal{X},\nu)$ on $0,\infty)$ such that $f_{0}=f$ and it holds
(2.4) $\frac{\mathrm{d}}{\mathrm{d}t}f_{t}\in-\partial^{-}\operatorname{Ch}(f_{t}),\quad\text{a.e. }t>0,$
which is the gradient flow of the Cheeger energy. The operator $P_{t}$ given by $f_{0}\mapsto f_{t}$ is known as the heat semigroup in $L^{2}(\mathcal{X},\nu)$.
In [6] properties of this gradient flow have been studied extensively, especially in connection with concepts coming from optimal transport theory like Wasserstein distance and relative entropy. In particular, the main results of [6] are critical in the identification of the gradient flows for the Cheeger energy in $L^{2}(\mathcal{X},\nu)$ and for the entropy functional in $(\mathscr{P}_{2}(\mathcal{X}),\mathsf{W}_{2})$.
@@
It is well known that Sobolev-type inequalities (including log-Sobolev inequalities) imply Talagrand’s transport inequalities, as shown by Otto and Villani [58, 73] on Euclidean spaces and Riemannian manifolds and extended by Gigli and Ledoux [33] to Polish metric measure spaces. A nonlinear analogue of this result was given by F. Y. Wang [75, Theorem 2.1] which is proven by extending Otto and Villani’s idea. Specifically, F. Y. Wang showed that if a weighted manifold $(\mathcal{M},\nu)$ admits a $\theta$-nonlinear log-Sobolev inequality, then the following nonlinear transport inequality holds:
@@
(2.5) $\mathsf{W}_{2}\left(\mu,\nu\right)\leq\Upsilon(0,D(\mu\|\nu)),\;\forall\mu\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{M}),$
where
$\Upsilon(s,\alpha):=\int_{s}^{\alpha}\frac{1}{\sqrt{\theta(r)}}{\text{\rm d}}r.$
In the present paper, we require a variant of F. Y. Wang’s result. Given $\mu_{W}$, we consider the conditional Wasserstein metric $\mathsf{W}_{2}\left(\mu_{X|W},\mu_{Y|W}|\mu_{W}\right)$ as a distance between conditional probabilities $\mu_{X|W}$ and $\mu_{Y|W}$.
@@
Let $(\mathcal{X},d)$ be a Polish (complete and separable) metric space and $\nu$ is a Borel probability measure on $\mathcal{X}$ which admits a $\theta$-nonlinear log-Sobolev inequality for some function $\theta$. Let $\mu_{X_{0}W}$ be a joint probability measure such that $\mu_{X_{0}}=f_{0}\nu\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{X})$ and $D(\mu_{X_{0}|W}\|\nu|\mu_{W})<\infty$. Let $\mu_{X_{t}}=f_{t}\nu$ be the heat flow with initial data $\mu_{X_{0}}$ and denote the induced transition probability $\mu_{X_{t}|X_{0}}$. Let $\mu_{X_{t}X_{0}W}=\mu_{X_{0}W}\mu_{X_{t}|X_{0}}$. Then, the curve $t\mapsto\mu_{X_{t}|W}$ is absolutely continuous on $0,\infty)$ with respect to the conditional Wasserstein metric $\mathsf{W}_{2}$ and for its metric speed $|\dot{\mu}_{X_{t}|W}|$, it holds that $|\dot{\mu}_{X_{t}|W}|^{2}\leq I(\mu_{X_{t}|W}\|\nu|\mu_{W})$ for a.e. $t$. Moreover, the map $t\mapsto D(\mu_{X_{t}|W}\|\nu|\mu_{W})$ is absolutely continuous and non-increasing on $0,\infty)$, and
(2.6) $\mathsf{W}_{2}\left(\mu_{X_{0}|W},\mu_{X_{t}|W}|\mu_{W}\right)\leq\Upsilon(D(\mu_{X_{t}|W}\|\nu|\mu_{W}),D(\mu_{X_{0}|W}\|\nu|\mu_{W})).$
Remark 5.
If $W$ is constant, then the unconditional version is obtained. This theorem can be generalized to the $\Phi$-entropy version by incorporating some ideas from [73] to the proof.
@@
Since $\mu_{X_{0}}\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{X})$, we have $\mu_{X_{0}|W=w}\in\mathscr{P}_{2}^{\mathrm{ac}}(\mathcal{X})$ for $\mu_{W}$-a.e. $w$. Observe that for $0\leq s<t\leq 1$,
$\mathsf{W}_{2}^{2}\left(\mu_{X_{s}|W},\mu_{X_{t}|W}|\mu_{W}\right)$ $=\mathbb{E}_{\mu_{W}}[\mathsf{W}_{2}^{2}\left(\mu_{X_{s}|W},\mu_{X_{t}|W}\right)]$
$\leq(t-s)\mathbb{E}_{\mu_{W}}[\int_{s}^{t}I(\mu_{X_{r}|W}\|\nu){\text{\rm d}}r]$
(2.7) $=(t-s)\int_{s}^{t}I(\mu_{X_{r}|W}\|\nu|\mu_{W}){\text{\rm d}}r,$
where the inequality follows since $\mathsf{W}_{2}\left(\mu_{X_{s}|W=w},\mu_{X_{t}|W=w}\right)\leq(t-s)\int_{s}^{t}I(\mu_{X_{r}|W=w}\|\nu){\text{\rm d}}r$ for $\mu_{W}$-a.e. $w$; see [6, p. 359]. Hence, for $0\leq s<t\leq 1$,
$\mathsf{W}_{2}\left(\mu_{X_{s}|W},\mu_{X_{t}|W}|\mu_{W}\right)\leq\frac{1}{2}(t-s+\int_{s}^{t}I(\mu_{X_{r}|W}\|\nu|\mu_{W}){\text{\rm d}}r).$
By the triangle inequality for $\mathsf{W}_{2},$ this inequality extends to the case of $0\leq s<t<\infty$. The integral here is finite for any finite $t$ since
$\int_{0}^{t}I(\mu_{X_{r}|W}\|\nu|\mu_{W}){\text{\rm d}}r$ $=\mathbb{E}_{\mu_{W}}[\int_{0}^{t}I(\mu_{X_{r}|W}\|\nu){\text{\rm d}}r]$
(2.8) $=\mathbb{E}_{\mu_{W}}[D(\mu_{X_{0}|W}\|\nu)-D(\mu_{X_{t}|W}\|\nu)]$
(2.9) $=D(\mu_{X_{0}|W}\|\nu|\mu_{W})-D(\mu_{X_{t}|W}\|\nu|\mu_{W}),$
and hence
(2.10) $\int_{0}^{\infty}I(\mu_{X_{r}|W}\|\nu|\mu_{W}){\text{\rm d}}r\leq D(\mu_{X_{0}|W}\|\nu|\mu_{W})<\infty$
where (2.8) follows by [6, Proposition 4.22]. Hence, $t\mapsto\mu_{X_{t}|W}$ is absolutely continuous on $0,\infty)$. From (2.7), we estimate $|\dot{\mu}_{X_{t}|W}|^{2}\leq I(\mu_{X_{t}|W}\|\nu|\mu_{W})$ for a.e. $t$.
Equations (2.9) and (2.10) imply that $t\mapsto D(\mu_{X_{t}|W}\|\nu|\mu_{W})$ is absolutely continuous and nonincreasing on $0,\infty)$. In fact, (2.9) further implies that $\frac{{\text{\rm d}}}{{\text{\rm d}}t}D(\mu_{X_{t}|W}\|\nu|\mu_{W})=-I(\mu_{X_{t}|W}\|\nu|\mu_{W})$ for a.e. $t$.
We next prove (2.6) by following Otto and Villani’s idea in [58] and Gigli and Ledoux’s extension [33]. For $\mu_{W}$-a.e. $w$, $\theta(D(\mu_{X|W=w}\|\nu))\leq I(\mu_{X|W=w}\|\nu).$ Taking expectation and by Jensen’s inequality,
$\theta(D(\mu_{X|W}\|\nu|\mu_{W}))\leq\mathbb{E}_{\mu_{W}}[\theta(D(\mu_{X|W}\|\nu))]\leq I(\mu_{X|W}\|\nu|\mu_{W}).$
Define
$f(t)=\int_{0}^{t}\sqrt{I(\mu_{X_{r}|W}\|\nu|\mu_{W})}{\text{\rm d}}r+\Upsilon(0,D(\mu_{X_{t}|W}\|\nu|\mu_{W})).$
Then, it holds that for a.e. $t$,
$\frac{{\text{\rm d}}^{+}}{{\text{\rm d}}t}f(t)$ $\leq\sqrt{I(\mu_{X_{t}|W}\|\nu|\mu_{W})}+\frac{\frac{{\text{\rm d}}}{{\text{\rm d}}t}D(\mu_{X_{t}|W}\|\nu|\mu_{W})}{\sqrt{\theta(D(\mu_{X_{t}|W}\|\nu|\mu_{W}))}}$
$=\sqrt{I(\mu_{X_{t}|W}\|\nu|\mu_{W})}-\frac{I(\mu_{X_{t}|W}\|\nu|\mu_{W})}{\sqrt{\theta(D(\mu_{X_{t}|W}\|\nu|\mu_{W}))}}\;\leq 0,$
where $\frac{{\text{\rm d}}^{+}}{{\text{\rm d}}t}$ is the right derivative.
So, $f$ is non-increasing as a function of $t$, which implies $f(t)\leq f(0)=\Upsilon(0,D(\mu_{X_{0}|W}\|\nu|\mu_{W}))$ for any $t\geq 0$. Thus,
$\int_{0}^{t}\sqrt{I(\mu_{X_{r}|W}\|\nu|\mu_{W})}{\text{\rm d}}r\leq\Upsilon(D(\mu_{X_{t}|W}\|\nu|\mu_{W}),D(\mu_{X_{0}|W}\|\nu|\mu_{W})).$
Combining it with $|\dot{\mu}_{X_{t}|W}|^{2}\leq I(\mu_{X_{t}|W}\|\nu|\mu_{W})$ for a.e. $t$. yields (2.6). ∎
We immediately obtain a corollary for Gaussian measure.